OpenStudy (anonymous):

if A/B= 3/4 then 3B=____

OpenStudy (anonymous):

12

OpenStudy (anonymous):

just plug it in

OpenStudy (anonymous):

a. 3/4 b. 4/3 c. 4a d. 3b theese are the answer options

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A = 3 and B = 4... so 3b would be 3/4

OpenStudy (mysesshou):

one of the options to pick is 3B=3b ? so 3b is a possible answer? I'm not sure it is 3/4 if B=4, then 3B = 3*4 = 12

OpenStudy (mysesshou):

probably